# VideoLifter: Lifting Videos to 3D with Fast Hierarchical Stereo Alignment

### v2

@@ Abstract
Efficiently reconstructing 3D scenes from monocular video remains a core challenge in computer vision, vital for applications in virtual reality, robotics, and scene understanding. Recently, frame-by-frame progressive reconstruction without camera poses is commonly adopted, incurring high computational overhead and compounding errors when scaling to longer videos. To overcome these issues, we introduce VideoLifter, a novel video-to-3D pipeline that leverages a local-to-global strategy on a fragment basis, achieving both extreme efficiency and SOTA quality. Locally, VideoLifter leverages learnable 3D priors to register fragments, extracting essential information for subsequent 3D Gaussian initialization with enforced inter-fragment consistency and optimized efficiency. Globally, it employs a tree-based hierarchical merging method with key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency while efficiently mitigating cumulative errors. This approach significantly accelerates the reconstruction process, reducing training time by over 82% while holding better visual quality than current SOTA methods.
@@ 1 Introduction
Efficiently reconstructing 3D scenes from sequential image observations is a longstanding challenge in computer vision, with applications spanning AR/VR, video processing, and autonomous driving. Recently, reconstructing 3D scenes from single video sequences (video-to-3D) has gained significant traction. This trend is driven by two factors: the increasing accessibility of handheld capture devices, making video capture more practical for non-professional users, and recent advancements in high-fidelity 3D reconstruction methods such as Neural Radiance Fields (NeRF) [30] and 3D Gaussian Splatting (3D-GS) [21].
Figure 1: Novel View Synthesis and Training Time Comparisons. VideoLifter does not require precomputed camera parameters (i.e., camera intrinsics K from COLMAP), reduces the training time required by the most relevant baseline CF-3DGS [14] by 82% while improving image quality (SSIM).
@@
Most video-to-3D reconstruction methods, based on either NeRF or 3D-GS, heavily depend on Structure-from-Motion (SfM) [36] to generate initial sparse reconstructions, providing essential components like camera poses, intrinsics, and an initial point cloud to build dense 3D models. However, when applied to video data, SfM is often unreliable or even infeasible (Issue ➊), because it relies on photometric assumptions that frequently break down in low-texture or challenging lighting conditions [28, 31, 36], though some works improve SfM for some specific conditions [25]. In response, recent methods [14, 6, 24, 46] have shifted toward jointly optimizing camera poses and scene representations rather than relying solely on SfM-based initializations. Nevertheless, these approaches still depend on accurate camera intrinsics from SfM, limiting their applicability in in-the-wild video scenarios.
@@
More importantly, those SfM-free video-to-3D methods typically reconstruct scenes incrementally from a canonical view, with two critical issues. First, they are slow and inefficient (Issue ➋) due to an iterative, frame-by-frame approach that re-optimizes the entire sequence with each new frame, thereby prolonging training times ($>$ 2 hours) and complicating the handling of complex trajectories, especially given the video setup, not a few images. Second, they are susceptible to incremental errors (Issue ➌), as the frame-by-frame approach tends to accumulate errors over long video sequences.
To address these issues, we propose VideoLifter, a novel video-to-3D reconstruction pipeline that achieves a 5$\times$ speed-up and enhanced view-synthesis quality compared to state-of-the-art methods, as demonstrated in Fig. 1.
@@
We effectively adopt the local-to-global stream to handle long-sequence videos on a fragment basis and then subsequently merge fragments into a final, globally consistent 3D scene. Our pipeline is driven by two key innovations that make the local-to-global concept workable with significantly boosted efficiency (Issue ➋) and much-reduced incremental errors(Issue ➌). First, in the Fragment Registration with Learned 3D Priors (Local) stage, we extract essential information (e.g., pointmaps and local camera poses for 3D Gaussian initialization) from each fragment by leveraging pretrained prior models such as MASt3R [23] to address Issue ➊. Rather than naively using 3D priors to initialize 3D Gaussians, like InstantSplat [12], we improve efficiency by (1) enforcing inter-fragment consistency via solely considering the key frames (set to first frame), solved on an efficient subgraph instead complete graph, and (2) extracting only the essential parameters (6-dimensional quaternion pose and 1-dimensional scale) for each view within fragment, thereby avoiding costly global optimization of full point maps. Second, in the Hierarchical Gaussian Alignment (Global) stage, we merge fragments through a tree-based hierarchical framework that employs key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency and mitigate cumulative errors efficiently.
Overall, our main contributions are as follows:
@@
  • Our fragment registration with learned 3D priors efficiently extracts essential representations for subsequent dense 3D-GS with several key efficiency-driven optimizations along with leverage learned priors to remove reliance on traditionally module SfM.
@@ 2 Related Works
3D reconstruction for high-quality novel view synthesis is the task of generating unseen views of a scene or object from a set of given images [1, 29]. NeRF [30] is one of the most influential methods, but it remains computationally expensive and slow. Subsequent work has focused on either improving the quality of renderings [2, 3, 4] or boosting the efficiency of the approach [34, 37, 41, 40], though few methods succeed in both quality and efficiency simultaneously due to the ill-posed nature of transforming a sparse point cloud into a dense 3D scene representation. Recently, unstructured radiance field methods have emerged [7, 21, 48], using different scene representation primitives. Among these, 3D-GS [21] stands out and has demonstrated impressive performance in efficiently reconstructing complex, real-world scenes with high fidelity. Both NeRFs and 3DGS rely on carefully captured sequential video or multi-view images to ensure sufficient scene coverage, utilizing preprocessing tools like SfM software, e.g., COLMAP [36], to compute camera parameters and provide a sparse SfM point cloud as additional input. Specifically, 3D-GS performs the sparse-to-dense optimization starting from SfM point clouds, named adaptive density control, critical for converting sparse point clouds into dense 3D Gaussian representations.
@@
Estimating 3D structure and camera motion simultaneously is a well-explored challenge [15, 33, 47]. SfM has seen significant advancements across various dimensions. Methods like [26, 17] focus on enhancing feature detection, while [39] introduces innovative optimization approaches. Improved data representations and more robust structural solutions are explored by [16, 36]. Simultaneous Localization and Mapping (SLAM) systems, on the other hand, track camera motion and construct 3D structures from video sequences [33, 11, 32]. Despite these advances, traditional SfM and SLAM techniques remain vulnerable to issues such as low-texture regions, occlusions, moving objects, and lighting variations, limiting their overall robustness and performance, though some works [25] aim to improve it. Recent papers [43, 9, 10, 38] finds using NN-based methods to estimate matching/pose estimation, which finds more robust than the hand-designed feature SIFT.
@@
Inaccuracies from SfM can propagate through subsequent radiance field reconstruction, reducing overall quality. Various approaches have been proposed to eliminate the reliance on SfM by jointly optimizing camera parameters and scene representation. For instance, NeRFmm [46] and BARF [24] optimize camera intrinsics/extrinsics alongside the radiance field, while SCNeRF [19] address camera distortions via specialized parameterizations. GARF [19] further simplifies the joint optimization and improve both efficiency and accuracy by using Gaussian-MLP models. SPARF [44] and TrackNeRF [27] introduces a method to simulate pose noise by injecting Gaussian noise into the camera parameters.
Recently, depth priors from monocular depth estimator are used to guide radiance fields optimization [6, 8, 28, 14]. More recent work, such as [20] and InstantSplat [12], either supplement depth priors with other priors (e.g., image matching network), or integrate end-to-end stereo models like DUST3R/MASt3R to reduce the dependency on camera pose information. While these methods show promise in removing the SfM reliance, scaling them effectively to a large number of views remains a challenge.
Figure 2: CF-3DGS’s frame-by-frame pipeline accumulates errors, causing long-term drift, while our method could largely alleviate it. Results are tested on 247_26441_50907 from CO3D-V2.
@@ 3.1 Video-to-3D: Challenges and Our Design
Video-to-3D Reconstruction:
Given a sequence of $N$ unposed and uncalibrated images from a monocular video, denoted as =$\mathcal{I}=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, VideoLifter aims to reconstruct the scene using 3D Gaussians $\mathcal{G}$ along with estimated camera intrinsics $K$ and extrinsics $\mathcal{T}=\{T_{i}\in\mathbb{R}^{3\times 4}\}_{i=1}^{N}$. We assume that all frames share a common intrinsic matrix, as they are from a single monocular video.
Key Challenges:
While NeRF and 3DGS have advanced 3D reconstruction, their variants remain suboptimal for video-to-3D reconstruction in terms of speed and quality. Existing methods often adopt a frame-by-frame progressive reconstruction paradigm, making them inherently slow and prone to cumulative errors when processing long sequence videos. As a result, long-term performance drift often occurs, as illustrated in Fig. 2. Furthermore, they typically rely on SfM to estimate camera intrinsic, which is unreliable or even infeasible for in-the-wild video sequences.
Our High-level Framework:
To meet the critical need for efficiency and quality in long-sequence video to 3D reconstruction, we depart from conventional frame-by-frame or holistic optimization methods by embracing a hierarchical local-to-global design philosophy. Specifically, we process long video sequences on a fragment basis and subsequently merge these fragments into a single, consistent 3D scene. Although the local-to-global concept has been explored in domains such as SLAM, adapting it to a video-to-3D reconstruction pipeline with 3DGS is new with two key, unanswered challenges:
@@
  • How can we merge multi-fragments into a high-quality and consistent 3D scene without alignment issues?
In response, we propose an efficient and high-quality video-to-3D reconstruction pipeline, VideoLifter, as illustrated in Fig.3. Our method achieves SOTA performance in both speed and quality (see Fig.1) through two key contributions: Local: Fragment Registration with Learned 3D Priors with efficiency-driven designs (Sec. 3.2) and Global: Hierarchical Gaussian Alignment to compress alignment error (Sec. 3.3).
@@ 3.2 Fragment Registration with Learned 3D Priors
Here, we describe our method to efficiently and reliably extract essential 3D reconstruction information on a fragment basis and the way we enforce cross-fragment consistency.
@@
Fragmentation Method Choice: In this work, we use a straightforward fragmentation strategy by uniformly dividing the frames into disjoint windows. Although simple, this approach has proven effective on benchmark datasets. More sophisticated methods, such as using frame-to-frame similarity to guide fragmentation, are orthogonal to our core contributions and could further enhance VideoLifter, e.g., in videos with abrupt view changes during capture.
Fragment-level Challenges: Naively applying existing methods for the fragment level (e.g., those in LocalRF [28] or CF-3DGS [14]) is neither efficient enough nor does it adequately prepare for future merging. Challenge ➀: SfM is heavily relied for NeRF/3D-GS, while it is not always available or reliable—especially in our video-to-3D reconstruction with varying conditions. Challenge ➁: A critical issue in the local-to-global paradigm is ensuring that fragments can be merged without incurring significant alignment errors. Challenge ➂: Beyond the inherent efficiency benefits of a local-to-global design, further enhancements in efficiency are necessary at the video setup.
@@
Inspired by recent work on replacing SIFT with NN-based method, e.g., LoFTR [43]/GIM [38], we abandon reliance on SfM and instead leverage learned 3D priors from large-scale pretrained foundation models. In this work, we use MASt3R [23] as prior model since it seamlessly integrates both geometric and matching cues. We emphasize that our video-to-3D pipeline are not exclusively tailored to MASt3R; rather, our VideoLifter is flexible and can incorporate any model that provides robust geometric and matching priors. For example, in Tab. 1, we demonstrate that VideoLifter performs well with both MASt3R and alternative approaches (e.g. LoFTR [43] for matching cues, Metric3Dv2 [18] for geometric cues). We choose MASt3R for its simplicity and efficiency.
@@
LoFTR [43] Metric3Dv2 [18] 0.9238 31.30 0.0757 0.005
@@
Key Frames as Anchor (Challenge ➁, ➂): In a fragment-based approach, ensuring inter-fragment consistency is critical for high-quality 3D reconstruction. To address this, we propose to “anchor” each fragment with a key frame, set to its first frame, $I_{(i-1)k+1}$ and enforce consistency only among these key frames. This strategy simplifies the consistency problem by limiting the problem scale in $\frac{N}{k}$ key frames instead of considering whole $N$ frames, thereby improve efficiency and reduce complexity.
@@
Key frame consistency is enforced at the point-cloud level by generating a globally optimized dense point map, along with transformation matrices across adjacent fragments $\{T^{f}_{i\rightarrow i+1}\}_{i=1}^{m-1}$. To achieve this, we follow a procedure similar to that in MASt3R [23], optimizing:
$\small(\tilde{\bm{P}}^{*},\bm{T}_{e}^{*})=\arg\min_{\tilde{\bm{P}},\bm{T},\sigma}\sum_{e\in\mathcal{E}}\sum_{v\in e}\sum_{i=1}^{HW}\bm{O}_{v,e}^{i}\left\|\tilde{\bm{P}}_{v}^{i}-\sigma_{e}\bm{T}_{e}\bm{P}_{v,e}^{i}\right\|.$
@@
where for each image pair $e=(v,u)\in\mathcal{E}$, $\sigma_{e}$ is scale factor, $\bm{P}_{v,e}$ and $\bm{O}_{v,e}$ is the pointmap and confidence map of $v$, respectively. However, MASt3R builds a complete graph for this optimization, resulting in a complexity of $\mathcal{O}((\frac{N}{k})^{2})$, it becomes inefficient enough for long video sequences.
@@
To enhance efficiency, we propose to build a more efficient sub-graph that only builds edges between the key frames and their four closest neighboring frames. This design is motivated by the observation that neighboring segments share greater co-visibility; hence, edges between key frames with distant neighbors can be safely pruned. This sub-graph greatly reduces the optimization complexity to $\mathcal{O}(4\frac{N}{k})$, which scales linearly with #frames $N$, while still showing high end-to-end 3D reconstruction quality.
@@
Efficient intra-Fragment Feature Registration (Challenge ➂): Finally, we aim to obtain an initial estimate of the local camera poses and pointmaps along with depth scale factors within each non-overlapping fragment, which can accelerate and boost the quality of the subsequent 3D Gaussian construction. A naive solution is to follow InstantSplat [12] to use MASt3R to solve the intra-fragment problem, but it requires optimizing millions of points and camera poses, making it inefficient.
@@
To enhance efficiency, we leverage the pre-obtained key frame information and obtain the needed information only considering pairwise relationships between the key frame and all subsequent frames in the same fragment. In this way, we only need to solve for 6-dimensional camera poses (in quaternion format) and a 1-dimensional scale factor for each view, which has proven sufficient to maintain end-to-end reconstruction quality at high efficiency. Moreover, we believe that this simplified, non-sequential matching approach can reduce the incremental errors that are commonly encountered in sequential matching.
@@
Camera pose: We refine the relative camera poses within the fragment using initial pairwise estimates. First, we identify the intersection of 2D correspondences between the key frame $I_{1}$ and each subsequent frame from index 2 to $k$. This process yields a consistent set of correspondences across all frames in the fragment. Using these intersected 2D correspondences, we retrieve the corresponding 3D positions from the key frame, which were previously optimized during key frame processing. These 3D-2D correspondences are then input to PnP-RANSAC [13], refining the camera poses to ensure alignment with consistent 3D points across views within the fragment. Only a 6-dim quaternion pose is optimized instead of directly optimizing pointmaps.
Scale factor: Scale variations may persist within the point clouds of the fragment due to independent inference. To address this, intersected 3D points from the key frame are utilized for scale estimation across all image pairs. Specifically, for each image pair $(I_{1},I_{i})$ between $I_{1}$ and $\{I_{i}\}_{i=2}^{k}$, the corresponding 3D point positions (with a total of $P$ points) are retrieved. A one-degree-of-freedom (1-DoF) scale factor is computed between the intersected 3D points in the current pair and those from the key frame:
@@
Time (min) Peak GPU Mem (GB) Time (min) Peak GPU Mem (GB)
@@
Table 2: Time and peak GPU memory usage on A6000 for varying view counts: MASt3R global optimization vs. our Sec. 3.2.
Overall Efficiency/Quality Improvement with Our Novel Fragment Registration: In contrast to naively applying MASt3R’s global optimization as InstantSplat [12] to compute point clouds and camera poses for all images, our method significantly enhances efficiency. As in Tab. 2, their global optimization approach runs out of memory beyond 64 views and incurs up to an 18$\times$ longer inference time. Moreover, Tab. 5 demonstrates that replacing our fragment registration with MASt3R initialization yields lower-quality results. This is primarily because it trys to solve a more complex global problem with less accurate pointmaps/poses outputs, whereas our approach eases optimization complexity by providing a more robust initialization for subsequent 3D-GS, delivering better quality.

### v3

# VideoLifter: Lifting Videos to 3D with Fast and Efficient Hierarchical Stereo Alignment

@@ Abstract
Efficiently reconstructing 3D scenes from monocular video remains a core challenge in computer vision, vital for applications in virtual reality, robotics, and scene understanding. Recently, frame-by-frame progressive reconstruction without camera poses is commonly adopted, incurring high computational overhead and compounding errors when scaling to longer videos. To overcome these issues, we introduce VideoLifter, a novel video-to-3D pipeline that leverages a local-to-global strategy on a fragment basis, achieving both extreme efficiency and SOTA quality. Locally, VideoLifter leverages learnable 3D priors to register fragments, extracting essential information for subsequent 3D Gaussian initialization with enforced inter-fragment consistency and optimized efficiency. Globally, it employs a tree-based hierarchical merging method with key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency while efficiently mitigating cumulative errors. This approach significantly accelerates the reconstruction process, reducing training time by over 82% while holding better visual quality than SOTA methods.
Figure 1: Novel View Synthesis and Training Time Comparisons. VideoLifter does not require precomputed camera parameters (i.e., camera intrinsics K from COLMAP), reduces the training time required by the most relevant baseline CF-3DGS fu2023colmap by 82% while improving image quality (SSIM).
@@ 1 Introduction
Reconstructing 3D scenes from image observations is a longstanding problem in computer vision, with applications spanning AR/VR, video processing, and autonomous driving. Recently, reconstructing 3D scenes from a single video (video-to-3D) has gained significant traction. This trend is driven by two factors: the increasing accessibility of handheld capture devices, making video capture more practical for non-professional users, and recent advancements in high-fidelity 3D reconstruction methods such as Neural Radiance Fields (NeRF) mildenhall2021nerf and 3D Gaussian Splatting (3D-GS) kerbl20233d .
@@
Most video-to-3D reconstruction methods heavily depend on Structure-from-Motion (SfM) schonberger2016structure to generate initial sparse reconstructions, providing essential components like camera poses, intrinsics, and the initial point cloud to build dense 3D models using NeRF or 3DGS. However, when applied to video data, SfM is often unreliable or even infeasible (Issue ➊), because it relies on photometric assumptions that frequently break down in low-texture or challenging lighting conditions meuleman2023progressively ; moulon2013global ; schonberger2016structure , though some works improve SfM for some specific conditions lindenberger2021pixel . In response, recent methods fu2023colmap ; bian2023nope ; lin2021barf ; wang2021nerf have shifted toward jointly optimizing camera poses and scene representations rather than relying solely on SfM-based initializations. But these approaches still depend on accurate camera intrinsics from SfM, limiting their applicability in in-the-wild video scenarios.
@@
More importantly, those SfM-free video-to-3D methods typically reconstruct scenes incrementally from a canonical view, with two critical issues. First, they are slow and inefficient (Issue ➋) due to an iterative, frame-by-frame approach that re-optimizes the entire sequence with each new frame, thereby prolonging training times ($>$ 2 hours) and complicating the handling of complex trajectories, especially given the video setup, not a few images. Naively using non-incremental InstantSplat fan2024instantsplat cannot handle video data with an Out-of-memory (OOM) issue, which cannot scale to many frames (See Tab. 2). Second, they are susceptible to incremental errors (Issue ➌), as the frame-by-frame approach tends to accumulate errors over long video sequences.
@@
To address these issues, we propose VideoLifter, a novel video-to-3D reconstruction pipeline that achieves a 5$\times$ speed-up and enhanced novel view-synthesis quality compared to state-of-the-art methods, as demonstrated in Fig. 1. We effectively adopt the local-to-global idea stream to handle long-sequence videos on a fragment basis and then subsequently merge fragments into a final, globally consistent 3D scene. Our pipeline is driven by two key innovations that make the local-to-global concept workable with significantly boosted efficiency (Issue ➋) and much-reduced incremental errors (Issue ➌) on video-to-3D. First, in the Fragment Registration with Learned 3D Priors (Local) stage, we extract essential information (e.g., pointmaps and local camera poses for 3D Gaussian initialization) from each fragment by leveraging pretrained prior models such as MASt3R leroy2024grounding to address Issue ➊. Rather than naively using 3D priors to initialize 3D Gaussians, like InstantSplat (MASt3R + 3D-GS) fan2024instantsplat , we improve efficiency by (1) enforcing inter-fragment consistency via solely considering the key frames, solved on an efficient subgraph instead complete graph, and (2) extracting only the essential parameters (6-dimensional quaternion pose and 1-dimensional scale) for each view within fragment, thereby avoiding costly global optimization of full point maps. Second, in the Hierarchical Gaussian Alignment (Global) stage, we merge fragments through a tree-based hierarchical framework that employs key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency and mitigate cumulative errors efficiently. Overall, our main contributions are as follows:
@@
  • Our fragment registration with learned 3D priors efficiently extracts essential representations for subsequent dense 3D-GS with several key efficiency-driven optimizations along with learned 3D priors to remove reliance on traditional module SfM.
@@ 2 Related Works
3D reconstruction for high-quality novel view synthesis generates unseen views of a scene or object from a set of images avidan1997novel ; mildenhall2019local . After the seminal NeRF work mildenhall2021nerf , a wave of unstructured radiance field methods has emerged kerbl20233d ; xu2022point , each adopting different scene-representation primitives. Among these, 3D-GS kerbl20233d stands out with impressive performance in efficiently reconstructing complex, real-world scenes with high fidelity. Both NeRFs and 3DGS rely on carefully captured sequential video or multi-view images to ensure sufficient scene coverage, utilizing preprocessing tools like SfM software, e.g., COLMAP schonberger2016structure , to compute camera parameters and provide a sparse SfM point cloud as additional input.
@@
Estimating 3D structure and camera motion is a well-explored challenge furukawa2010towards ; newcombe2011dtam ; wu2011visualsfm . SfM has seen significant advancements across various dimensions. Methods like lowe2004distinctive ; MatchNet focus on enhancing feature detection, snavely2011scene introduces innovative optimization methods, gherardi2010improving ; schonberger2016structure explore improved data representations and more robust structural solutions. Despite these advances, traditional SfM techniques remain vulnerable to issues such as low-texture regions, occlusions, moving objects, and lighting variations, limiting their overall robustness and performance.
@@
Inaccuracies from SfM can propagate through subsequent radiance field reconstruction, reducing overall quality. Various approaches have been proposed to eliminate the reliance on SfM by jointly optimizing camera parameters and scene representation, such as NeRFmm wang2021nerf and BARF lin2021barf . GARF jeong2021self further simplifies the joint optimization and improve both efficiency and accuracy by using Gaussian-MLP models. SPARF truong2023sparf and TrackNeRF mai2024tracknerf introduces a method to simulate pose noise by injecting Gaussian noise into the camera parameters. Recently, depth priors from monocular depth estimator are used to guide radiance fields optimization bian2023nope ; cheng2023lu ; meuleman2023progressively ; fu2023colmap . More recent work, such as jiang2024construct and InstantSplat fan2024instantsplat , either supplement depth priors with other priors (e.g., image matching network), or integrate end-to-end stereo models like DUST3R/MASt3R to reduce the dependency on camera pose information. While these methods show promise in removing the SfM reliance, scaling them to a large number of views remains a challenge, such as the OOM issue (e.g., InstantSplat), drifting error (e.g., CF-3DGS), and unsatisfactory quality (e.g., InstantSplat).
Comparison with Simultaneous Localization and Mapping (SLAM)
Although both SLAM and video-to-3D reconstruction process multiple views, their input conditions and end goals differ fundamentally. First, SLAM operates online, processing frames sequentially as they arrive, whereas video-to-3D has access to the entire sequence upfront. This offline setting enables pipelines to consider all frames together rather than a purely sequential approach. Second, our primary objective is novel view synthesis, generating photorealistic views from unseen viewpoints, which SLAM methods are not designed to support, but only to rerender the training set sandstrom2024splat . Hence, SLAM-based techniques are not directly applicable to the video-to-3D reconstruction problem.
@@ 3.1 Video-to-3D: Challenges and Our Design
Video-to-3D Reconstruction
Given a sequence of $N$ unposed and uncalibrated images from a monocular video, denoted as =$\mathcal{I}=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, VideoLifter aims to reconstruct the scene
Figure 2: CF-3DGS’s frame-by-frame pipeline accumulates errors with long-term drift, while our method compresses drift error along # frames. Results are tested on 247_26441_50907 from CO3D-V2.
using 3D Gaussians $\mathcal{G}$ along with estimated camera intrinsics $K$ and extrinsics $\mathcal{T}=\{T_{i}\in\mathbb{R}^{3\times 4}\}_{i=1}^{N}$. We assume that all frames share a common intrinsic matrix, as they are from a single monocular video.
Key Challenges
While NeRF and 3D-GS have advanced 3D reconstruction, their variants remain suboptimal for video-to-3D reconstruction in terms of speed and quality. Existing methods often adopt a frame-by-frame progressive reconstruction method, making them inherently slow and prone to cumulative errors (See long-term performance drift in Fig. 2) when processing videos. Furthermore, they typically rely on SfM to estimate camera intrinsic, which is unreliable or even infeasible for in-the-wild video sequences.
Our High-level Framework
To meet the critical need for efficiency and quality in long-sequence video to 3D reconstruction, we depart from conventional frame-by-frame or holistic optimization methods by embracing a hierarchical local-to-global design philosophy. Specifically, we process long video sequences on a fragment basis and subsequently merge these fragments into a single, consistent 3D scene. Although the local-to-global concept is not new, adapting it to a video-to-3D reconstruction pipeline with 3DGS is new, with two key, unanswered challenges:
@@
How can we merge fragments into a high-quality and consistent 3D scene without alignment issues?
In response, we propose VideoLifter, an efficient and high-quality video-to-3D reconstruction pipeline, as illustrated in Fig. 3. Our method achieves SOTA performance in both speed and quality (see Fig. 1) through two-level innovations: Local: Fragment Registration with Learned 3D Priors with efficiency-driven designs (Sec. 3.2) and Global: Hierarchical Gaussian Alignment to compress alignment error (Sec. 3.3).
@@ 3.2 Fragment Registration with Learned 3D Priors
First, we describe our method to efficiently and reliably extract essential 3D reconstruction information on a fragment basis and the way we enforce cross-fragment consistency.
@@
Fragmentation Method Choice: In this work, we use a straightforward fragmentation strategy by uniformly dividing the frames into disjoint windows. Although simple, this approach has proven effective on benchmarks. More sophisticated methods, such as using frame-to-frame similarity to guide fragmentation, could further enhance VideoLifter, e.g., in videos with abrupt view changes during capture, but are orthogonal to our core contributions.
Fragment-level Challenges: Naively applying existing methods for the fragment level (e.g., those in LocalRF meuleman2023progressively or CF-3DGS fu2023colmap ) is neither efficient enough nor does it adequately prepare for future merging. Challenge ➀: SfM is heavily relied for NeRF/3D-GS, while it is not always available or reliable—especially in our video-to-3D reconstruction with varying conditions. Challenge ➁: A critical issue in the local-to-global paradigm is ensuring that fragments can be merged without incurring significant alignment errors. Challenge ➂: Beyond the inherent efficiency benefits of a local-to-global design, further enhancements in efficiency are necessary at the video setup.
@@
Inspired by recent work on replacing SIFT with NN-based method, e.g., LoFTR sun2021loftr /GIM shen2024gim ,
@@
LoFTR sun2021loftr Metric3Dv2 hu2024metric3d 0.9238 31.30 0.0757 0.005
@@
we abandon reliance on SfM and instead leverage learned 3D priors from large-scale pretrained foundation models. In this work, we use MASt3R leroy2024grounding as prior model since it seamlessly integrates both geometric and matching cues. We emphasize that our video-to-3D pipeline are not exclusively tailored to MASt3R; rather, our VideoLifter is flexible and can incorporate any model that provides robust geometric and matching priors. For example, in Tab. 1, we demonstrate that VideoLifter performs well with both MASt3R and alternative approaches (e.g. LoFTR sun2021loftr for matching cues, Metric3Dv2 hu2024metric3d for geometric cues). We choose MASt3R for its simplicity and efficiency.
@@
Key Frames as Anchor (Challenge ➁, ➂): In a fragment-based approach, ensuring inter-fragment consistency is critical for high-quality 3D reconstruction. To address this, we propose to “anchor" each fragment with a key frame, set to its first frame, $I_{(i-1)k+1}$ and enforce consistency only among these key frames. This strategy simplifies the consistency problem by limiting the problem scale in $\frac{N}{k}$ key frames instead of considering whole $N$ frames, thereby improve efficiency and reduce complexity.
@@
Key frame consistency is enforced at the point-cloud level by generating a globally optimized dense point map, along with transformation matrices across adjacent fragments $\{T^{f}_{i\rightarrow i+1}\}_{i=1}^{m-1}$. To achieve this, we follow a procedure similar to that in MASt3R leroy2024grounding , optimizing:
$\small(\tilde{\bm{P}}^{*},\bm{T}_{e}^{*})=\arg\min_{\tilde{\bm{P}},\bm{T},\sigma}\sum_{e\in\mathcal{E}}\sum_{v\in e}\sum_{i=1}^{HW}\bm{O}_{v,e}^{i}\left\|\tilde{\bm{P}}_{v}^{i}-\sigma_{e}\bm{T}_{e}\bm{P}_{v,e}^{i}\right\|.$ (1)
@@
where for each image pair $e=(v,u)\in\mathcal{E}$, $\sigma_{e}$ is scale factor, $\bm{P}_{v,e}$ and $\bm{O}_{v,e}$ is the pointmap and confidence map of $v$, respectively. However, MASt3R builds a complete graph for this optimization, resulting in a complexity of $\mathcal{O}((\frac{N}{k})^{2})$, which becomes inefficient enough for long video sequences.
@@
To enhance efficiency rather than naively using MASt3R-like method, we propose to build a more efficient sub-graph that only builds edges between the key frames and their four closest neighboring frames. This design is motivated by the observation that neighboring segments share greater co-visibility; hence, edges between key frames with distant neighbors can be safely pruned. This sub-graph greatly reduces the optimization complexity to $\mathcal{O}(4\frac{N}{k})$, which scales linearly with #frames $N$, while still showing high end-to-end 3D reconstruction quality.
@@
Efficient intra-Fragment Feature Registration (Challenge ➂): Finally, we aim to obtain an initial estimate of the local camera poses and pointmaps along with depth scale factors within each non-overlapping fragment, which can accelerate and boost the quality of the subsequent 3D Gaussian construction. A naive solution is to follow InstantSplat fan2024instantsplat to use MASt3R to solve the intra-fragment problem, but it requires optimizing millions of points and camera poses, making it inefficient.
@@
To enhance efficiency, we use the pre-obtained key frame information and obtain the needed information only considering pairwise relationships between key frame and all subsequent frames in the same fragment. In this way, we only need to solve for 6-dimensional camera poses (in quaternion format) and a 1-dimensional scale factor for each view. We found it sufficient to maintain end-to-end reconstruction quality with high efficiency. Moreover, this simplified, non-sequential matching approach can reduce the incremental errors that are commonly encountered in sequential matching(See Fig. 2).
@@
Camera pose: We refine the relative camera poses within the fragment using initial pairwise estimates. First, we identify the intersection of 2D correspondences between the key frame $I_{1}$ and each subsequent frame from index 2 to $k$. This process yields a consistent set of correspondences across all frames in the fragment. Using these intersected 2D correspondences, we retrieve the corresponding 3D positions from the key frame, which were previously optimized during key frame processing. These 3D-2D correspondences are then input to PnP-RANSAC fischler1981random , refining the camera poses to ensure alignment with consistent 3D points across views within the fragment. Only a 6-dim quaternion pose is optimized instead of directly optimizing pointmaps.
Scale factor: Scale variations may persist within the point clouds of the fragment due to independent inference. To address this, intersected 3D points from the key frame are utilized for scale estimation across all image pairs. Specifically, for each image pair between $I_{1}$ and $\{I_{i}\}_{i=2}^{k}$, the corresponding 3D point positions (with a total of $P$ points) are retrieved. A one-degree-of-freedom (1-DoF) scale factor is computed between the intersected 3D points in the current pair and those from the key frame:
@@
Overall Efficiency/Quality Improvement with Our Novel Fragment Registration: In contrast
@@
Time
Peak GPU
Mem (GB)
Time
Peak GPU
Mem (GB)
@@
Table 2: Time and peak GPU memory usage on an A6000 for varying view counts: Direct using MASt3R global optimization vs. our Sec. 3.2.
to naively applying MASt3R’s global optimization as InstantSplat fan2024instantsplat to compute point clouds and camera poses for all images, our method significantly enhances efficiency. As in Tab. 2, their global optimization approach runs out of memory beyond 64 views and incurs up to an 18$\times$ longer inference time. Moreover, Tab. 5 demonstrates that replacing our fragment registration with MASt3R initialization yields lower-quality results. This is primarily because it trys to solve a more complex global problem with less accurate pointmaps/poses outputs, whereas our approach eases optimization complexity by providing a more robust initialization for subsequent 3D-GS, delivering better quality.